# DECODE: Domain-aware Continual Domain Expansion for Motion Prediction

### v1

@@ V-B Main Results
Table I presents the performance evaluation of our proposed DECODE framework by comparing performance metrics across several learning paradigms. For each column, we illustrate how the performance evolves after each learning phase, denoted as $\tt{minADE}_{\tt{Di},\tt{Pj}}$ and $\tt{minFDE}_{\tt{Di},\tt{Pj}}$ where $j\geq i$ indicates the metrics for each domain $i$ after learning phase $j$. For consistency, all models utilize the same MTR model structure. This lineup includes the original MTR model denoted as oriMTR, which is trained solely on current domain datasets without any continual learning strategies. For comparative purposes, we also include the pre-trained generalized MTR model trained on the WOMD, denoted as preMTR, which undergoes no additional training on new domains. Additionally, we assess versions of MTR that integrate an online Elastic Weight Consolidation (EWC) strategy (ewcMTR), and another that employs an experience replay buffer (erMTR).
The results highlight significant performance fluctuations in the oriMTR, which lacks any continual learning strategy. In contrast, the performance of the preMTR remains static, as it does not engage in any continual learning. Notably, the generalized model’s performance, while stable, does not surpass that of oriMTR when initially trained on a domain. A substantial increase in displacement errors is evident when transitioning from the RounD to the HighD domain, attributable to the distinct driving patterns of these domains. However, performance improves upon moving to the InD domain, reflecting the similarity in driving conditions between RounD and InD. Interestingly, the performance of ewcMTR is significantly inferior to that of erMTR, which indicates that experience replay may be more effective than EWC in this context. Overall, the DECODE strategy demonstrates the most stable performance compared to other methods. It exhibits only a slight increase in errors during the initial learning phase of each domain and maintains relatively stable error rates thereafter.
@@
Figure 6: Visualization of qualitative performance across different learning phases in the DECODE framework: Each row represents repeated observations of the same scene to highlight the consistency and variation in the model’s predictive behavior as it transitions through phases. Columns delineate the different learning phases, demonstrating how predictions evolve with the progression of training and adaptation to new information.
@@
Following the procedure outlined in Algorithm 1, the most suitable specialized model is selected based on the highest log evidence generated by the normalizing flow models. Table III presents a confusion matrix for the RounD, HighD, and InD domains, illustrating the classification performance across different validation sets. The data in the table indicate that the majority of scenes are correctly classified, with the predictions being generated by the specialized model that is most familiar with each respective scene. The overall classification accuracy is exceptionally high, with an accuracy of 0.998, a precision of 0.998, and a recall of 0.997, demonstrating the effectiveness of our model selection process based on domain familiarity.
TABLE III: Confusion Matrix of Three Learning Phases
@@
Initially, training commenced with the normal domain and subsequently expanded to include the conflict domain. Table IV presents the confusion matrix for the normalizing flow models, demonstrating an accuracy of 0.986, a precision of 0.974, and a recall of 0.985. These results highlight the model’s capability to effectively distinguish between different vehicular behaviors, which is essential for providing crucial safety information in the planning task.
TABLE IV: Confusion Matrix of Normal vs. Conflict Domains
@@
TABLE V: Comparison between models with and without deep Bayesian uncertainty estimation
@@
($\tt{minADE}$/$\tt{minFDE}$) ($\tt{minADE}$/$\tt{minFDE}$) ($\tt{minADE}$/$\tt{minFDE}$)
@@ V-C Ablation Study
To evaluate the impact of dynamic balancing between generalized and specialized models using deep Bayesian uncertainty estimation, Table V presents a comparison of our proposed model’s performance with and without this feature. We analyze the model after phase 1, which includes only one specialized model trained on the RounD domain, assessing its performance on the HighD and InD domains. For the dynamic balancing variant, we set the prior evidence at $10$. The results indicate improved performance in both domains when dynamic balancing is implemented, with particularly notable enhancements observed in the HighD domain. This improvement substantiates the necessity of dynamic balancing for the framework to manage unfamiliar scenarios effectively. Further analysis of the sensitivity of various prior evidence settings is depicted in Fig. 9. Increasing prior evidence generally enhances performance in both the HighD and InD domains until it stabilizes. The HighD domain exhibits a significant initial decrease in minADE and minFDE. However, as prior evidence continues to increase, the performance in the RounD domain deteriorates, indicating an over-reliance on the generalized model even though the specialized model shows superior results.
@@ VI Conclusion and Future Work
Looking ahead, we aim to further enhance the DECODE framework by incorporating foundation models to provide a deeper, more generalized understanding of various driving scenarios. This would involve refining our approach to domain definition and moving beyond geographic and situational specifics to encompass more complex and nuanced settings. Additionally, we plan to explore the architectural optimization of hypernetworks, considering designs where parameters remain fixed, and domain expansion is driven solely by new domain queries. These advancements will focus on broadening the adaptability and precision of our framework.

### v2

@@ V-B Main Results
Table I presents the performance evaluation of our proposed DECODE framework by comparing performance metrics across multiple learning phases. For each column, we illustrate how performance evolves after each learning stage, denoted as $\tt{minADE}_{\tt{Di},\tt{Pj}}$ and $\tt{minFDE}_{\tt{Di},\tt{Pj}}$, where $j\geq i$ indicates the metric for domain $i$ after learning phase $j$. For consistency, all models utilize the same MTR backbone architecture. The comparison lineup includes: (1) the original MTR model (oriMTR)[62], trained from scratch on each current domain dataset without any continual learning strategy; (2) a pre-trained generalized MTR model (preMTR) trained on the whole WOMD dataset, used directly without further domain adaptation; (3) regularization-based continual learning baselines, including Elastic Weight Consolidation (EWC) [38] and Memory Aware Synapses (MAS) [2]; (4) rehearsal-based baselines, including a standard Experience Replay (ER) buffer and Dark Experience Replay (DER++) [6]; and (5) state-of-the-art parameter-efficient tuning methods, namely DualPrompt [70] and CODA-Prompt [65], adapted to operate with the pre-trained MTR backbone.
The results highlight significant performance fluctuations in the oriMTR, which lacks any continual learning strategy. In contrast, the performance of the preMTR remains static, as it does not engage in any continual learning. Notably, the generalized model’s performance, while stable, does not surpass that of oriMTR when initially trained on a domain. A substantial increase in displacement errors is evident when transitioning from the RounD to the HighD domain, attributable to the distinct driving patterns of these domains. However, performance improves upon moving to the InD domain, reflecting the similarity in driving conditions between RounD and InD. Regularization-based methods like EWC and MAS perform notably worse than rehearsal-based approaches such as ER and DER++, indicating that replaying past samples is more effective than parameter regularization for reducing forgetting in motion prediction. Noticeably, since the original DER++ framework applies soft-label matching on logits for pure classification tasks, in our motion prediction setting we adopt this strategy only for modality updates while using standard rehearsal for regression updates. For both DualPrompt and CODA-P, we report results using the prompts inserted into both encoder and decoder, and our proposed method consistently outperforms or matches these approaches across all learning phases. Overall, the DECODE strategy demonstrates the most stable performance among all compared methods, exhibiting only a slight increase in errors during the initial phase of each new domain and maintaining relatively steady performance thereafter.
@@
Figure 6: Visualization of qualitative performance across different learning phases in the DECODE framework: Each row represents repeated observations of the same scene to highlight the consistency and variation in the model’s predictive behavior as it transitions through phases. Columns delineate the different learning phases, demonstrating how predictions evolve with the progression of training and adaptation to new information. In each subfigure, yellow bars represent the log evidence computed by the normalizing flow for each specialized model, while the posterior weights assigned to the generalized (G) and specialized (S) models are shown below.
@@
As shown in Table II, our DECODE framework achieves performance comparable to or better than SOTA PET–based continual learning methods. Both DualPrompt and CODA-P were originally designed for image classification with a ViT-B/16 [17] backbone that uses only an encoder with self-attention layers, To ensure fair comparison, we adapted their prompting designs to fit the encoder–decoder MTR structure: for DualPrompt, we followed the original G-Prompt and E-Prompt configurations—placing G-Prompts in the first two encoder layers and E-Prompts in the last four—and added a variant extending E-Prompts to all six decoder layers; for CODA-P, we tested encoder-only, decoder-only, and encoder–decoder variants with prompt components inserted at corresponding stages. Both methods used the same prompt hyperparameters reported in their original papers. As summarized in Table III, our DECODE framework consistently outperforms or matches them across all settings, demonstrating superior adaptability and effectiveness for continual learning in motion prediction.
TABLE III: Performance comparison of PET variants
Method $\tt{AER}\downarrow$ $\tt{FGT}\downarrow$ ($\tt{minADE}$ / $\tt{minFDE}$) ($\tt{minADE}$ / $\tt{minFDE}$) Dualprompt enc $0.794\pm 0.010$ / $2.053\pm 0.029$ $054\pm 0.006 / 0.0$$027\pm 0.012$ enc+dec $0.608\pm 0.009$ / $1.509\pm 0.037$ $\mathbf{0.040\pm 0.002}$ / $0.062\pm 0.009$ Coda-P enc $0.721\pm 0.007$ / $1.871\pm 0.015$ $0.255\pm 0.008$ / $0.541\pm 0.021$ dec $0.647\pm 0.006$ / $1.633\pm 0.018$ $0.280\pm 0.006$ / $0.698\pm 0.017$ enc+dec $0.685\pm 0.006$ / $1.728\pm 0.017$ $0.415\pm 0.007$ / $1.020\pm 0.016$ DECODE $\mathbf{0.584\pm 0.005}/\mathbf{1.395\pm 0.019}$ $0.044\pm 0.001/\mathbf{0.027\pm 0.001}$
@@
TABLE IV: Confusion Matrix of Three Learning Phases
@@
TABLE V: Confusion Matrix of Normal vs. Conflict Domains
@@
Following the procedure outlined in Algorithm 1, the most suitable specialized model is selected based on the highest log evidence generated by the normalizing flow models. Table IV presents a confusion matrix for the RounD, HighD, and InD domains, illustrating the classification performance across different validation sets. The data in the table indicate that the majority of scenes are correctly classified, with the predictions being generated by the specialized model that is most familiar with each respective scene. The overall classification accuracy is exceptionally high, with an accuracy of 0.998, a precision of 0.998, and a recall of 0.997, demonstrating the effectiveness of our model selection process based on domain familiarity.
Initially, training commenced with the normal domain and subsequently expanded to include the conflict domain. Table V presents the confusion matrix for the normalizing flow models, demonstrating an accuracy of 0.986, a precision of 0.974, and a recall of 0.985. These results highlight the model’s capability to effectively distinguish between different vehicular behaviors, which is essential for providing crucial safety information in the planning task.
TABLE VI: Comparison between models with and without deep Bayesian uncertainty estimation
@@
($\tt{minADE}$$\tt{minFDE}$) ($\tt{minADE}$$\tt{minFDE}$) ($\tt{minADE}$$\tt{minFDE}$)
@@ V-C Ablation Study
To evaluate the impact of dynamic balancing between generalized and specialized models using deep Bayesian uncertainty estimation, Table VI presents a comparison of our proposed model’s performance with and without this feature. We analyze the model after phase 1, which includes only one specialized model trained on the RounD domain, assessing its performance on the HighD and InD domains. For the dynamic balancing variant, we set the prior evidence at $10$. The results indicate improved performance in both domains when dynamic balancing is implemented, with particularly notable enhancements observed in the HighD domain. This improvement substantiates the necessity of dynamic balancing for the framework to manage unfamiliar scenarios effectively. Further analysis of the sensitivity of various prior evidence settings is depicted in Fig. 9. Increasing prior evidence generally enhances performance in both the HighD and InD domains until it stabilizes. The HighD domain exhibits a significant initial decrease in minADE and minFDE. However, as prior evidence continues to increase, the performance in the RounD domain deteriorates, indicating an over-reliance on the generalized model even though the specialized model shows superior results. In practice, we follow a simple hyperparameter tuning strategy to choose the value of the prior evidence. Specifically, we incrementally increase $e^{(0)}$ starting from zero and select the value that yields the best performance on the main evaluation metric (minADE) over a small buffer of representative samples from each domain, assuming access to limited validation data. This approach provides a practical and effective guideline for setting the prior evidence in a data-driven manner.
@@
V-D Computational and Parameter Efficiency Analysis
We further compare the parameter and runtime efficiency. As shown in Fig. 10a, the parameter overhead and growth pattern across learning phases are illustrated. We compare DECODE with all DualPrompt and CODA-P variants. The rest methods maintain the same parameter count as the original MTR model. Our framework introduces the largest initial parameter overhead due to the inclusion of a hypernetwork, but its growth across learning phases is negligible since only a small domain query component is added per new task. In contrast, although DualPrompt begins with minimal overhead, it requires progressively larger prompt sets as new tasks are introduced. CODA-P maintains a fixed prompt pool and exhibits a parameter count comparable to DECODE.
Figure 10: Parameter and runtime efficiency comparison among continual learning methods across different learning phases. (a) Parameter growth of each method, measured as the overhead relative to the baseline model (MTR). (b) Inference latency (line plot) and GPU memory footprint (circle size), illustrating the runtime overhead of each method.
Fig. 10b presents the inference latency per frame and GPU memory usage across learning phases. For a fair comparison, we evaluate on the rounD domain with up to 8 target vehicles per frame. While DECODE initially exhibits similar latency to other methods, its computation time grows linearly and faster as the number of domains increases. This is primarily due to the need to evaluate the normalizing flow for each known specialized domain during inference.
To address this limitation, we highlight two future optimization directions:
  • Hypernetwork optimization — our current implementation performs model-wise looping instead of batching across domains, which could be significantly accelerated.
  • Domain pruning — since domain shifts between consecutive frames are typically small, the set of candidate domains can be constrained based on recent similarity metrics. Additionally, organizing domains into a hierarchical structure could make best-domain selection more scalable as the number of specialized models grows.
V-E Extension to General Continual Learning Tasks
TABLE VII: Results on 5-datasets (10 classes per task)
Method $\tt{Avg.ACC}\uparrow$ $\tt{FGT}\downarrow$
ER(500) $84.26\pm 0.84^{\dagger}$ $12.85\pm 0.62^{\dagger}$
DER++(500) $84.88\pm 0.57^{\dagger}$ $10.46\pm 1.02^{\dagger}$
EWC $50.93\pm 0.09^{\dagger}$ $34.94\pm 0.07^{\dagger}$
LwF $47.91\pm 0.33^{\dagger}$ $38.01\pm 0.28^{\dagger}$
F2P $81.14\pm 0.93^{\dagger}$ $4.64\pm 0.52^{\dagger}$
Dualprompt $88.08\pm 0.36^{\dagger}$ $2.21\pm 0.69^{\dagger}$
Coda-P $84.82\pm 0.45^{\ddagger}$ $13.24\pm 0.45^{\ddagger}$
DECODE $\mathbf{90.56\pm 1.43}$ $\mathbf{0.05\pm 0.01}$
[…]
TABLE VIII: Results on 10-task CIFAR-100 (10 classes per task)
Method $\tt{Avg.ACC}\uparrow$ $\tt{FGT}\downarrow$
ER(5000) $82.53\pm 0.17^{\dagger}$ $16.46\pm 0.25^{\dagger}$
DER++(5000) $83.94\pm 0.34^{\dagger}$ $14.55\pm 0.73^{\dagger}$
EWC $47.01\pm 0.29^{\dagger}$ $33.27\pm 1.17^{\dagger}$
LwF $60.69\pm 0.63^{\dagger}$ $27.77\pm 2.17^{\dagger}$
F2P $83.86\pm 0.28^{\dagger}$ $7.35\pm 0.38^{\dagger}$
Dualprompt $86.51\pm 0.33^{\dagger}$ $5.16\pm 0.09^{\dagger}$
Coda-P $\mathbf{91.71\pm 0.78}^{\ddagger}$ $\mathbf{3.26\pm 0.83}^{\ddagger}$
DECODE $86.41\pm 1.23$ $4.55\pm 0.51$
[…]
The DECODE framework is not limited to motion prediction and can be extended to general continual learning applications, such as image classification, which is widely studied in the literature. To demonstrate this, we adapt DECODE to work with a ViT-B/16 model [17] pre-trained on the ImageNet-1k [58] dataset, using it as the backbone and generalized model. Unlike motion prediction models that often adopt an encoder–decoder architecture, the ViT-B/16 model consists of an encoder followed by a classification head. To integrate DECODE, we make the following modifications:
We consider the encoder with $K=12$ self-attention layers, and define two parallel computational routes. The first route passes the final ($K$-th) layer’s output to produce a [CLS] token, which is then used as input to the normalizing flow for domain inference. The second route branches from the (K1)(K{-}1)-th layer and connects to a specialized self-attention layer followed by a specialized classification head, whose parameters are generated by the hypernetwork. This design preserves the core structure of the original ViT model while enabling domain specialization, demonstrating the flexibility of our framework to support task-specific adaptation across diverse continual learning scenarios.
Table VII presents results on a five-dataset benchmark [18] composed of diverse image classification tasks: CIFAR-10[42], MNIST[43], SVHN[52], notMNIST[5], and Fashion-MNIST[74]. This setting reflects a realistic continual learning scenario with high task heterogeneity. Table VIII reports results on the widely used 10-task CIFAR-100 benchmark [42], a standard for evaluating class-incremental learning. During training, we freeze the pretrained ViT backbone in the same way as the prompt-based baselines used for comparison, and only train the hypernetwork. All methods are trained with the same data exposure and number of optimization steps (25 epochs for five-dataset and 40 epochs for CIFAR-100), ensuring comparable training budgets. For each experiment, we conduct five training trials and report the average accuracy and forgetting, following the definitions in [70]. Across both benchmarks, our DECODE framework consistently achieves competitive or superior performance compared to state-of-the-art continual learning methods. Notably, DECODE performs particularly well on the five-dataset benchmark, where we observe that the normalizing flow is more effective when task diversity is high.
@@ VI Conclusion and Future Work
Looking ahead, we plan to enhance DECODE by optimizing the hypernetwork architecture—keeping parameters fixed while expanding through new domain queries. We also aim to improve the efficiency of the domain awareness module by batching operations and bounding the number of active domains per frame to reduce latency. These efforts will further broaden DECODE’s adaptability and precision.